OpenStudy (nthenic_oftime):

Begin by graphing the standard absolute value function f(x) = | x |. Then use transformations of this graph to describe the graph the given function. h(x) = 2 | x | + 2

OpenStudy (anonymous):

I don't remember how to do this sorry

OpenStudy (anonymous):

f(x)=|x| g(x)=f(x) + 1 g(x)=|x| + 1 h(x)=2 g(x) h(x) = 2(f(x) + 1) h(x) = 2|x| + 2

OpenStudy (nthenic_oftime):

@OllieT are all those after the original the descriptions of the transformations or did you just give me a step?

OpenStudy (nthenic_oftime):

@saxprincess325 thats okay :) thank you

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

I created a g(x) so then it would show you how I got to h(x)

OpenStudy (anonymous):

@nthenic_oftime do you think you can help me on mine

OpenStudy (nthenic_oftime):

okay so that will suffice as my answer? thank you @OllieT

OpenStudy (nthenic_oftime):

tag me in it ill do my bhest @saxprincess325

OpenStudy (anonymous):

okayyyy thank you

OpenStudy (anonymous):

I would think so, what level is this? i.e GCSE AS level, A2?

OpenStudy (nthenic_oftime):

its in Pre-cal... but idk how accurately assigned the ciriculum is

OpenStudy (anonymous):

What Pre-Cal, sorry I'm from the UK

OpenStudy (nthenic_oftime):

pre calculus

OpenStudy (anonymous):

well I gathered that haha, I think it would be fine as an answer

OpenStudy (nthenic_oftime):

oh im sorry i didnt understand what yuou meant lol. but thank you :)